# Living at the Edge: A Critical Look at the Cosmological Neutrino Mass Bound

### v1

@@ IV.3 Planck + DESI + SN: Impact of the Dark Energy equation of state
Figure 6: Neutrino mass profile likelihoods for Planck+DESI-Y1+Pantheon+ data set combinations. We show $\Lambda$CDM in black, varying the equation of state of dark energy in green, and allowing for $A_{\rm lens}$ to vary in blue. In the left panel we show the results for plik, in the middle for CamSpec and in the right panel for Hillipop. We clearly see a similar behaviour for all of them and the potential preference for a negative best fit to dissapear when the equation of state of dark energy is allowed to vary.
@@ IV.4 Frequentists vs Bayesian Limits: The impact of statistical choices on the neutrino mass bound
Until now, we have focused our attention on profile likelihoods, as these allowed us to investigate the preference for negative neutrino masses. It remains to be seen however, how the confidence intervals built from the profile likelihood compares with the Bayesian credible intervals built from the posteriors, in order to understand how important is the choice of statistical procedure and to quantify the role of prior effect in the bound given by Eq. (1). We refer to Appendix C for the posterior distributions.
$\sum m_{\nu}$ $<0.084\,{\rm eV}\,\,\,[{\rm Bayesian}]\,,$ (5a)
$\sum m_{\nu}$ $<0.074\,{\rm eV}\,\,\,[{\rm Bounded\!-\!Likelihood}]\,,$ (5b)
$\sum m_{\nu}$ $<0.071\,{\rm eV}\,\,\,[{\rm Feldman\!-\!Cousins}]\,.$ (5c)
So far we have included in our analyses value for neutrino masses down to the massless limit, $\sum m_{\nu}=0$, but we know from the laboratory that there are physical boundaries at either $\sum m_{\nu}=0.06\,{\rm eV}$ for NO or at $\sum m_{\nu}=0.10\,{\rm eV}$ for IO. To gauge the impact of those experimental lower limits on the cosmological neutrino mass bound, we run dedicated Bayesian analyses restricting the prior to $\sum m_{\nu}$ following either the NO or IO constraints. For the frequentist limit, it is sufficient to consider these boundaries as lower limits in our $\Delta\chi^{2}$ curves. This procedure yields:
$\sum m_{\nu}$ $<0.121\,{\rm eV}\,\,\,[{\rm NO\!-\!Bayesian}]\,,$ (6a)
$\sum m_{\nu}$ $<0.106\,{\rm eV}\,\,\,[{\rm NO\!-\!Bounded\!-\!Likelihood}]\,,$ (6b)
$\sum m_{\nu}$ $<0.096\,{\rm eV}\,\,\,[{\rm NO\!-\!Feldman\!-\!Cousins}]\,,$ (6c)
and for the inverted ordering case:
$\sum m_{\nu}$ $<0.152\,{\rm eV}\,\,\,[{\rm IO\!-\!Bayesian}]\,,$ (7a)
$\sum m_{\nu}$ $<0.138\,{\rm eV}\,\,\,[{\rm IO\!-\!Bounded\!-\!Likelihood}]\,,$ (7b)
$\sum m_{\nu}$ $<0.127\,{\rm eV}\,\,\,[{\rm IO\!-\!Feldman\!-\!Cousins}]\,.$ (7c)
$\sum m_{\nu}$ $<0.125\,{\rm eV}\,\,\,[{\rm Bayesian}]\,,$ (8a)
$\sum m_{\nu}$ $<0.114\,{\rm eV}\,\,\,[{\rm Bounded\!-\!Likelihood}]\,,$ (8b)
$\sum m_{\nu}$ $<0.114\,{\rm eV}\,\,\,[{\rm Feldman\!-\!Cousins}]\,.$ (8c)
@@ V Do cosmological data prefer a “negative” neutrino mass?
Figure 7: Profile likelihoods for the data set combinations of Planck18-PR3+DESI-Y1 (black), HiLLiPoP23-PR4 (blue), Planck18-PR3+DESI-Y1-no07 (red), HiLLiPoP23-PR4+DESI-Y1-no07 (green), and compared with $\chi^{2}_{\rm eff}=-2\log\mathcal{P}$ from Green and Meyers (2024) (purple) and Elbers et al. (2024) (orange), which have treatments for “negative” neutrino masses. By comparing the black and blue curves we can clearly see that the bound on the neutrino mass gets relaxed if the HiLLiPoP likelihood (which does not contain a lensing anomaly) is used. However, it is clear from the extrapolated parabolas that there is still some preference for a negative neutrino mass. This, however, disappears when the two outliers at $z=0.7$ in DESI data are removed (see red and green curves).
@@
Our results agree overall fairly well with all previous works Craig et al. (2024); Allali and Notari (2024); Green and Meyers (2024); Elbers et al. (2024) considering the different treatments performed in each of them. Our main results and the comparison to previous analyses are shown in Fig. 7. In particular, Ref. Elbers et al. (2024) argued that the Gaussian extrapolation in Allali and Notari (2024) underestimates the preference for negative masses of present data. Indeed, Refs. Craig et al. (2024); Green and Meyers (2024); Elbers et al. (2024) find rather more negative best fits through their analyses with their respective “effective masses” peaking around $-0.15\,{\rm eV}$ (see purple and orange lines in Fig. 7). We do find some preference for negative neutrino masses in datasets including both the Planck likelihoods affected by the lensing anomaly (Plik 2018 in Fig. 7 but also CamSpec in previous sections) and the full DESI Y1 BAO data (black line). However, this preference is significantly weaker, with a best fit at $\sum m_{\nu}\simeq-0.05\,{\rm eV}$ in agreement with the extrapolation of the posterior shown in Allali and Notari (2024) and the comparison and discussion in Elbers et al. (2024).
@@
More importantly, Ref. Elbers et al. (2024) only considered the Plik 2018 Planck likelihood. We find here that this dataset is key in driving the preference for negative neutrino masses, because of the presence of a residual lensing anomaly, and that it can be reduced when using alternative likelihoods. Indeed, we find a preference for negative neutrino masses when Planck 2018 is analyzed on its own. Switching for the HiLLiPoP23-PR4 version of Planck (largely unaffected by the lensing anomaly), there is no preference for negative masses, and the bound significantly relax. The preference re-appears for HiLLiPoP23-PR4 when DESI-Y1 BAO data are included because these data have a lower constraining power to $\sum m_{\nu}$. Nevertheless, Fig. 7 shows that the bound on $\sum m_{\nu}$ relaxes when changing from plik 2018 to HiLLiPoP (black vs blue lines) or when removing the $z=0.7$ outlier bin in DESI (black vs red lines). Finally, when the outliers bin is removed and the HiLLiPoP implementation are both considered, the preference for negative masses essentially disappears (green line) and the bound is about a factor 2 weaker. Let us note that the authors of Green and Meyers (2024) did investigate the persistence of their effective negative neutrino mass signal when PR4 likelihoods are used. However, the analyses were not performed for HiLLiPoP23-PR4 but rather for the CamSpec22-PR4 likelihood, where the lensing anomaly is still present, albeit at a reduced level. Ref. Green and Meyers (2024) does show that the preference for negative masses is relaxed in this scenario at a level that, naively, seems consistent with the reduced presence of the anomaly in CamSpec. In this context, it would be very interesting to explore the potential preference for a “negative” neutrino mass following the implementation of Elbers et al. (2024) and Craig et al. (2024); Green and Meyers (2024) but using the HiLLiPoP likelihood and also in the absence of the $z=0.7$ outliers that we have shown pull the neutrino mass bound significantly. Given our results, and if our hypotheses are correct, the preference for negative neutrino masses would vanish in that case.
@@
From our results, we conclude that there is no significant nor compelling evidence for negative neutrino masses from cosmology. We do confirm some mild preference when Planck likelihoods affected by the lensing anomaly are adopted, pointing towards a potential residual systematic effect in these implementations. We also find that the $z=0.7$ bins of DESI, in some tension with Planck data, push towards negative values. If these outliers are confirmed, it will be interesting to see how the preference for negative masses evolves in the future. For now, and in agreement with Allali and Notari (2024), we conclude there is no compelling evidence for negative masses from present datasets and that present constraints are still perfectly compatible with the results from neutrino oscillations. We believe that a key check to see if cosmological data does prefer effective “negative” neutrino masses is to do a cosmological analysis using the HiLLiPoP likelihood implementation.
@@ VI Conclusions
  • Bayesian and frequentist limits on the neutrino mass agree very well and within 10% precision – The agreement between the Bayesian and frequentist bounds on $\sum m_{\nu}$ is very good. Regarding the frequentist constraints, an extrapolation into the unphysical region of negative neutrino masses was performed to apply the appropriate correction Feldman and Cousins (1998) to the confidence intervals. Even though this procedure introduces a certain degree of uncertainty, we quantified the range of variation with the number of points for larger values of $\sum m_{\nu}$ and $\chi^{2}$ included and verified that the impact on the confidence intervals derived is negligible. Moreover, we also compared these with the naive cuts at the values of $\Delta\chi^{2}$ implied by Wilks theorem without venturing into the unphysical region and found almost identical results. When comparing these frequentist confidence intervals with the Bayesian credible regions we find good agreement between both approaches but consistently tighter results for the frequentist method, by about $\sim 10\%$. We thus conclude that the cosmological constraints on $\sum m_{\nu}$ are robust at this level against variations of the statistical method employed to derive them.
@@
  • DESI-Y1 BAO measurements at $\boldsymbol{z=0.7}$ pull the neutrino mass bound significantly and also induce some preference for negative values – When combining Planck with BAO data from DESI we observe a mild preference for unphysical negative neutrino masses, and hence a more stringent constraint. This mild preference for negative masses even appears for the HiLLiPoP implementation, for which it is absent when analyzed alone or in combination with other datasets. Upon closer examination, this preference seems to be driven mostly by the two radial and angular BAO measurements at $z=0.7$, which are in $\sim\!3\sigma$ tension with Planck expectations. Indeed, when removing these outliers, the preference for negative masses disappears and the corresponding constraints relax accordingly. Similarly, if SDSS BAO data is used instead of DESI, constraints relax by about a $50\%$ in all cases considered. Our findings indicate that a statistical fluctuation or some systematic effect with the DESI measurements at $z=0.7$ could be behind the strong bound on the neutrino mass.
@@ Acknowledgments.
EFM acknowledges very interesting discussions with Stefano Gariazzo and thanks him for pointing out relevant references. This project has received support from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 860881-HIDDeN and No 101086085 - ASYMMETRY, and from the Spanish Research Agency (Agencia Estatal de Investigación) through the Grant IFT Centro de Excelencia Severo Ochoa No CEX2020-001007-S and Grant PID2022 137127NB-I00 funded by MCIN/AEI/10.13039/501100011033. We acknowledge support from the HPC-Hydra cluster at IFT and the HTCondor cluster at CERN. XM acknowledges funding from the European Union’s Horizon Europe Programme under the Marie Skłodowska-Curie grant agreement no. 101066105-PheNUmenal. The work of DNT was supported by the Spanish MIU through the National Program FPU (grant number FPU20/05333).
@@ Appendix B Comparison with other gaussian extrapolations
The comparison is shown in Fig. 11, where we plot our $\Delta\chi^{2}$ profiles and their $-2\log\mathcal{P}$ for similar datasets. We find an overall agreement in the preference for a best fit in the negative neutrino mass region, suggesting that the conclusions drawn from the extrapolation are not largely dependent on the statistical procedure employed

### v2

@@ IV.3 Planck + DESI + SN: Impact of the Dark Energy equation of state
Figure 8: Neutrino mass profile likelihoods for Planck+DESI-Y1+Pantheon+ data set combinations. We show $\Lambda$CDM in black, varying the equation of state of dark energy in green, and allowing for $A_{\rm lens}$ to vary in blue. In the left panel we show the results for plik, in the middle for CamSpec, and in the right panel for Hillipop. We clearly see a similar behaviour for all of them and the potential preference for a negative best fit to dissapear when the equation of state of dark energy is allowed to vary.
@@ IV.4 Frequentists vs Bayesian Limits: The impact of statistical choices on the neutrino mass bound
Until now, we have focused our attention on profile likelihoods, as these allowed us to investigate the preference for negative neutrino masses. In this section we will study how the confidence intervals built from the profile likelihood compare with the Bayesian credible intervals built from the posteriors. While the two statistical approaches address distinct questions and thus do not need to necessarily coincide, we find the comparison useful to understand how dependent is the derived constraint on the choice of the statistical procedure. Moreover, if the likelihood is Gaussian and when assuming flat, non-informative priors on the parameter under study, the two approaches should coincide. Therefore, the comparison allows to asses the role of possible prior effects in the bound given by Eq. (1) as well as deviations from Gaussianity. Notice that the extrapolations to negative masses performed in our work as well as in Ref. Alam et al. (2021) and Allali and Notari (2024) rely on the Gaussianity of the likelihood and posterior respectively. Similarly, the study of Ref. Elbers et al. (2024) relies on a linear expansion of the dependence of the observables with $\sum m_{\nu}$ which, per Wilks theorem, would imply a Gaussian behaviour. The fact that these analyses do not coincide (see section V for a full comparison) implies that the Gaussian behaviour approximation is violated at some degree. As we will show and discuss below, we find that the differences between the Bayesian and frequentist approaches are at the $10\%$ level, which provides a useful estimation of the size of the uncertainties introduced by these approximations. We refer to Appendix A for the posterior distributions.
@@
So far we have included in our analyses values for neutrino masses down to the massless limit, $\sum m_{\nu}=0$, but we know from the laboratory that there are physical boundaries at either $\sum m_{\nu}=0.06\,{\rm eV}$ for NO or at $\sum m_{\nu}=0.10\,{\rm eV}$ for IO. To gauge the impact of those experimental lower limits on the cosmological neutrino mass bound, we run dedicated Bayesian analyses restricting the prior to $\sum m_{\nu}$ following either the NO or IO constraints. For the frequentist limit, it is sufficient to consider these boundaries as lower limits in our $\Delta\chi^{2}$ curves. This procedure yields:
$\sum m_{\nu}$ $<0.121\,{\rm eV}\,\,\,[{\rm NO\!-\!Bayesian}]\,,$ (6a)
$\sum m_{\nu}$ $<0.106\,{\rm eV}\,\,\,[{\rm NO\!-\!Bounded\!-\!Likelihood}]\,,$ (6b)
$\sum m_{\nu}$ $<0.096\,{\rm eV}\,\,\,[{\rm NO\!-\!Feldman\!-\!Cousins}]\,,$ (6c)
and for the inverted ordering case:
$\sum m_{\nu}$ $<0.152\,{\rm eV}\,\,\,[{\rm IO\!-\!Bayesian}]\,,$ (7a)
$\sum m_{\nu}$ $<0.138\,{\rm eV}\,\,\,[{\rm IO\!-\!Bounded\!-\!Likelihood}]\,,$ (7b)
$\sum m_{\nu}$ $<0.127\,{\rm eV}\,\,\,[{\rm IO\!-\!Feldman\!-\!Cousins}]\,.$ (7c)
$\sum m_{\nu}$ $<0.125\,{\rm eV}\,\,\,[{\rm Bayesian}]\,,$ (8a)
$\sum m_{\nu}$ $<0.114\,{\rm eV}\,\,\,[{\rm Bounded\!-\!Likelihood}]\,,$ (8b)
$\sum m_{\nu}$ $<0.114\,{\rm eV}\,\,\,[{\rm Feldman\!-\!Cousins}]\,.$ (8c)
When considering the physical boundary for normal ordering, the constraints read:
$\sum m_{\nu}$ $<0.160\,{\rm eV}\,\,\,[{\rm NO-Bayesian}]\,,$ (9a)
$\sum m_{\nu}$ $<0.132\,{\rm eV}\,\,\,[{\rm NO-Bounded\!-\!Likelihood}]\,,$ (9b)
$\sum m_{\nu}$ $<0.125\,{\rm eV}\,\,\,[{\rm NO-Feldman\!-\!Cousins}]\,.$ (9c)
and for the inverted ordering case:
$\sum m_{\nu}$ $<0.179\,{\rm eV}\,\,\,[{\rm IO-Bayesian}]\,,$ (10a)
$\sum m_{\nu}$ $<0.156\,{\rm eV}\,\,\,[{\rm IO-Bounded\!-\!Likelihood}]\,,$ (10b)
$\sum m_{\nu}$ $<0.146\,{\rm eV}\,\,\,[{\rm IO-Feldman\!-\!Cousins}]\,.$ (10c)
Where, as before, the FC corrections become more relevant given the larger distance between the best fit and the physical boundary.
@@ V Do cosmological data prefer a “negative” neutrino mass?
Figure 9: Profile likelihoods for the data set combinations of Planck18-PR3+DESI-Y1 (black), HiLLiPoP23-PR4 (blue), Planck18-PR3+DESI-Y1-no07 (red), HiLLiPoP23-PR4+DESI-Y1-no07 (green), and compared with $\chi^{2}_{\rm eff}=-2\log\mathcal{P}$ from Green and Meyers (2024) (purple) and Elbers et al. (2024) (orange), which have treatments for “negative” neutrino masses. By comparing the black and blue curves we can clearly see that the bound on the neutrino mass gets relaxed if the HiLLiPoP likelihood (which does not contain a lensing anomaly) is used. However, it is clear from the extrapolated parabolas that there is still some preference for a negative neutrino mass. This, however, disappears when the DESI BAO data at $z=0.7$ which contains a $\sim 3\sigma$ outlier is removed (see red and green curves).
@@
Our results agree overall fairly well with all previous works Craig et al. (2024); Allali and Notari (2024); Green and Meyers (2024); Elbers et al. (2024) considering the different treatments performed in each of them. Our main results and the comparison to previous analyses are shown in Fig. 9 (see also App. B). In particular, Ref. Elbers et al. (2024) argued that the Gaussian extrapolation in Allali and Notari (2024) underestimates the preference for negative masses of present data. Indeed, Refs. Craig et al. (2024); Green and Meyers (2024); Elbers et al. (2024) find rather more negative best fits through their analyses with their respective “effective masses” peaking around $-0.15\,{\rm eV}$ (see purple and orange lines in Fig. 9). We do find some preference for negative neutrino masses in datasets including both the Planck likelihoods affected by the lensing anomaly (Plik 2018 in Fig. 9 but also CamSpec in previous sections) and the full DESI Y1 BAO data (black line). However, this preference is significantly weaker, with a best fit at $\sum m_{\nu}\simeq-0.05\,{\rm eV}$ in agreement with the extrapolation of the posterior shown in Allali and Notari (2024) and the comparison and discussion in Elbers et al. (2024).
@@
More importantly, Ref. Elbers et al. (2024) only considered the CamSpec Planck likelihood. We find here that this dataset has a preference for negative neutrino masses, because of the presence of a residual lensing anomaly, and that it can be reduced when using alternative likelihoods. Indeed, we find a preference for negative neutrino masses when Planck 2018 is analyzed on its own. Switching for the HiLLiPoP23-PR4 version of Planck (largely unaffected by the lensing anomaly), there is no preference for negative masses, and the bound significantly relaxes. The preference re-appears for HiLLiPoP23-PR4 when DESI-Y1 BAO data are included because these data have a lower constraining power to $\sum m_{\nu}$. Nevertheless, Fig. 9 shows that the bound on $\sum m_{\nu}$ relaxes when changing from plik 2018 to HiLLiPoP (black vs blue lines) or when removing the $z=0.7$ outlier bin in DESI (black vs red lines). Finally, when the $z=0.7$ bin is removed and the HiLLiPoP implementation are both considered, the preference for negative masses essentially disappears (green line) and the bound is about a factor 2 weaker. Let us note that the authors of Green and Meyers (2024) did investigate the persistence of their effective negative neutrino mass signal when PR4 likelihoods are used. However, the analyses were not performed for HiLLiPoP23-PR4 but rather for the CamSpec22-PR4 likelihood, where the lensing anomaly is still present, albeit at a reduced level. Ref. Green and Meyers (2024) does show that the preference for negative masses is relaxed in this scenario at a level that, naively, seems consistent with the reduced presence of the anomaly in CamSpec. In this context, it would be very interesting to explore the potential preference for a “negative” neutrino mass following the implementation of Elbers et al. (2024) and Craig et al. (2024); Green and Meyers (2024) but using the HiLLiPoP likelihood and also in the absence of the $z=0.7$ outliers that we have shown pull the neutrino mass bound significantly. Given our results, and if our hypotheses are correct, the preference for negative neutrino masses would vanish in that case.
@@
From our results, we conclude that there is no significant nor compelling evidence for negative neutrino masses from cosmology. We do confirm some mild preference when Planck likelihoods affected by the lensing anomaly are adopted, pointing towards a potential residual systematic effect in these implementations. We also find that the $z=0.7$ bin of DESI, in some tension with Planck data, push towards negative values. If these outliers are confirmed, it will be interesting to see how the preference for negative masses evolves in the future. For now, and in agreement with Allali and Notari (2024), we conclude there is no compelling evidence for negative masses from present datasets and that present constraints are still perfectly compatible with the results from neutrino oscillations. We believe that a key check to see if cosmological data does prefer effective “negative” neutrino masses is to do a cosmological analysis using the HiLLiPoP likelihood implementation.
@@ VI Conclusions
  • Bayesian and frequentist limits on the neutrino mass agree very well and within 10% precision – The agreement between the Bayesian and frequentist bounds on $\sum m_{\nu}$ is very good. Regarding the frequentist constraints, an extrapolation into the unphysical region of negative neutrino masses was performed to apply the appropriate correction Feldman and Cousins (1998) to the confidence intervals. Even though this procedure introduces a certain degree of uncertainty, we quantified the range of variation with the number of points for larger values of $\sum m_{\nu}$ and $\chi^{2}$ included and verified that the impact on the confidence intervals derived is negligible. Moreover, we also compared these with the naive cuts at the values of $\Delta\chi^{2}$ implied by Wilks theorem without venturing into the unphysical region and found almost identical results. When comparing these frequentist confidence intervals with the Bayesian credible regions with flat priors on $\sum m_{\nu}$ we find good agreement between both approaches but consistently tighter results for the frequentist method, by about $\sim 10\%$. We thus conclude that the cosmological constraints on $\sum m_{\nu}$ are robust at this level against variations of the statistical method employed to derive them.
@@
  • DESI-Y1 BAO measurements at $\boldsymbol{z=0.7}$ pull the neutrino mass bound significantly and also induce some preference for negative values – When combining Planck with BAO data from DESI we observe a mild preference for unphysical negative neutrino masses, and hence a more stringent constraint. This mild preference for negative masses even appears for the HiLLiPoP implementation, for which it is absent when analyzed alone or in combination with other datasets. Upon closer examination, this preference seems to be driven mostly the BAO measurements at $z=0.7$, out of which the angular one is in $\sim\!3\sigma$ tension with Planck expectations. Indeed, when removing this bin, the preference for negative masses disappears and the corresponding constraints relax accordingly. Similarly, if SDSS BAO data is used instead of DESI, constraints relax by about a $50\%$ in all cases considered. Thus, given the unexpectedly strong constraint derived when including the full DESI dataset, it will be very interesting to confirm if the present trend is confirmed with higher statistics from upcoming DESI data releases.
@@ Acknowledgments.
EFM acknowledges very interesting discussions with Stefano Gariazzo and thanks him for pointing out relevant references. We also warmly thank Thomas Schwetz for very interesting comments and suggestions for our analysis. This project has received support from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 860881-HIDDeN and No 101086085 - ASYMMETRY, and from the Spanish Research Agency (Agencia Estatal de Investigación) through the Grant IFT Centro de Excelencia Severo Ochoa No CEX2020-001007-S and Grant PID2022 137127NB-I00 funded by MCIN/AEI/10.13039/501100011033. We acknowledge support from the HPC-Hydra cluster at IFT and the HTCondor cluster at CERN. XM acknowledges funding from the European Union’s Horizon Europe Programme under the Marie Skłodowska-Curie grant agreement no. 101066105-PheNUmenal. The work of DNT was supported by the Spanish MIU through the National Program FPU (grant number FPU20/05333).
@@ Appendix B Comparison with other Gaussian extrapolations
The comparison is shown in Fig. 14, where we plot our $\Delta\chi^{2}$ profiles and their $-2\log\mathcal{P}$ for similar datasets. We find an overall agreement in the preference for a best fit in the negative neutrino mass region, suggesting that the conclusions drawn from the extrapolation are not largely dependent on the statistical procedure employed.